Table Manufacturing Company produce one style of tables. The following data pertain to producing one table.

Planned production per month:

Units (one table) = 100

Piece of Woods = 20

Estimate M price = $30

Actual production

Quantity purchased (QP) = 18

Actual Price (AP) = $34

The calculated material quantity (volume) variance is:

a. 72 Favorable

b. 72 Unfavorable

c. 60 Favorable

d. 60 Unfavorable

e. 12 Favorable

f. 12 Unfavorable
Answer #1

Materials Quantity Variance

= (SQ-AQ) * SP

= (20 - 18)* 30

= 2 * 30

= 60 Favourable

Option C

Comment if you face any issues